OpenStudy (anonymous):

Solve the following system. x ^2 + y ^2 = 20 y = x ^2 {(4, -2),(4, 2)} {(-2, 4), (2, 4)} (2, 4) (3, 9)

OpenStudy (shubhamsrg):

substitute x^2 =y in first equation. What do you get ?

OpenStudy (anonymous):

the answer i got was 2,4

OpenStudy (shubhamsrg):

(2,4) is absolutely correct. You just missed a root when you took square root on both sides. Note that if a = b^2 , then b = sqrt(a) and -sqrt(a). Now see if you can use this information to find the other root as well.